# Unfolding AlphaFold’s Bayesian Roots in Probability Kinematics

### v2

@@ 2.4 Probability kinematics
As in the classic example of Whitworth’s horses [16] (see Appendix 9), new information often does not arrive in the form of a deterministic event, but rather as a revision in the probability of various hypotheses. PK provides a framework for updating beliefs under such uncertain evidence. Unlike classical Bayesian conditioning, which updates a prior $\pi(\boldsymbol{\omega})$ to a posterior $p(\boldsymbol{\omega})\propto p(\boldsymbol{y}\mid\boldsymbol{\omega})\pi(\boldsymbol{\omega})$ where $\boldsymbol{y}$ is observed data, PK updating uses revised probabilities on a partition of $\boldsymbol{\omega}$ (Figure 1).
@@
when the conditional probabilities remain unchanged, i.e., $p(\boldsymbol{\omega}\mid\xi_{i})=\pi(\boldsymbol{\omega}\mid\xi_{i})$. This condition is referred to as the Jeffrey condition or J-condition.
Unlike conventional Bayesian inference, PK does not rely on an explicit likelihood function, and its evidence is not a concrete event. Nonetheless, PK preserves essential features of Bayesian updating. First, PK satisfies the Dutch book argument [3]. Second, PK reduces to conventional Bayesian updating when for one of the partition elements, $p(\xi_{i})=1$. Third, among all distributions that match the updated marginals $\{p(\xi_{i})\}$, the PK update minimizes the Kullback-Leibler divergence to the prior $\pi$ [16]. This links the method to a variational principle [7, 36] that fits the Bayesian interpretation of probability [16]. PK is usually formulated in terms of probabilities and finite partitions [48]. In the next section, we outline the more general probabilistic framework needed to address the AF1 case.
@@ 3.1 PK for the case of PDFs and infinite partitions
We provide a simplified definition tailored to the present case based on [16] and [48], while avoiding the formalism of measure theory and irrelevant boundary cases.
@@
  • A random variable $\boldsymbol{\omega}\in\Omega$, where $\Omega$ is a smooth manifold (e.g., the $m$-torus $\mathbb{T}^{m}$), equipped with a prior density $\pi(\boldsymbol{\omega})$.
  • A random variable $\boldsymbol{\xi}$, where $\boldsymbol{\xi}$ can be written as a many-to-one function $\boldsymbol{\xi}(\boldsymbol{\omega})$ of $\boldsymbol{\omega}$, taking values in a smooth manifold $\Xi$ (e.g., the space of valid pairwise distances between $n$ points in $\mathbb{R}^{k}$), forming an infinite partition of $\Omega$. Its prior density $\pi(\boldsymbol{\xi})$ is induced by $\pi(\boldsymbol{\omega})$ via
    $\pi(\boldsymbol{\xi})=\int\delta(\boldsymbol{\xi}(\boldsymbol{\omega})-\boldsymbol{\xi})\pi(\boldsymbol{\omega})\textrm{d}\boldsymbol{\omega}.$ (10)
    Each value $\boldsymbol{\xi}\in\Xi$ defines a partition element: the set $\{\boldsymbol{\omega}\in\Omega:\boldsymbol{\xi}(\boldsymbol{\omega})=\boldsymbol{\xi}\}$.
  • Evidence as a new density $p(\boldsymbol{\xi})$ (e.g., regarding pairwise distances between amino acids).
@@
    Update Rule: The posterior $p(\boldsymbol{\omega})$ is obtained as,
    $p(\boldsymbol{\omega})=\int\pi(\boldsymbol{\omega}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}=\pi(\boldsymbol{\omega}\mid\boldsymbol{\xi}(\boldsymbol{\omega}))p(\boldsymbol{\xi}(\boldsymbol{\omega})),.$ (13)
@@
    Reference Ratio ($R$) Update Rule: Equivalently, after applying Bayes’ formula to (13),
    $p(\boldsymbol{\omega})=\int\frac{\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))\pi(\boldsymbol{\omega})}{\pi(\boldsymbol{\xi})}p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}=\frac{p(\boldsymbol{\xi}(\boldsymbol{\omega}))}{\pi(\boldsymbol{\xi}(\boldsymbol{\omega}))}\pi(\boldsymbol{\omega}).$ (14)
    The Dirac delta function arises from the fact that $p(\boldsymbol{\xi}\mid\boldsymbol{\omega})=\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))$. Note that (14) follows directly from the J-condition given by (12).
@@ 3.2 KBPs versus PMFs
KBPs can be understood as an application of PK [25, 47, 67, 27]. Consider a prior distribution over protein structure, $\pi(\boldsymbol{X})$. Without any loss of generality, we leave out the conditioning on sequence. We are now given soft evidence about pairwise distances between amino acids, $p(\boldsymbol{D})$, say, only involving $C\beta$ atoms. Thus, the collective random variable $\boldsymbol{D}$ pertains to a subset of atoms, induces a partition on the space of $\boldsymbol{X}$, and corresponds to $\boldsymbol{\xi}$ in Definition 1. By construction, the J-condition holds, that is, $p(\boldsymbol{X}\mid\boldsymbol{D})=\pi(\boldsymbol{X}\mid\boldsymbol{D})$. In other words, if we sample an infinite number of structures from the prior and retain those with a specified $\boldsymbol{D}$, we would get samples from the desired distribution $p(\boldsymbol{X}\mid\boldsymbol{D})$. The PK update following (14) is then given by
$p(\boldsymbol{X})=\int\frac{\delta(\boldsymbol{D}(\boldsymbol{X})-\boldsymbol{D})\pi(\boldsymbol{X})}{\pi(\boldsymbol{D})}p(\boldsymbol{D})\textrm{d}\boldsymbol{D}=\frac{p(\boldsymbol{D}(\boldsymbol{X}))}{\pi(\boldsymbol{D}(\boldsymbol{X}))}\pi(\boldsymbol{X}).$ (15)
Note that $\boldsymbol{D}$ is a random variable, while $\boldsymbol{D}(\boldsymbol{X})$ is its realisation at $\boldsymbol{X}$. A minus-log transform recovers the KBP given by (8),
@@
if we assume conditional independence of pairwise distances (see also Appendix 11). The additional second term does not appear in standard KBP expressions, as it is typically implicitly accounted for by Monte Carlo sampling [57] or scoring of existing structures [60].
The contrast of KBPs with PMFs is now obvious: while PMFs amount to a projection and marginalization of a well-defined Boltzmann distribution for a single physical system as in (6), KBPs as in (16) correspond to a generalized Bayesian update of an (empirical) prior by means of PK. PK and PMFs produce distributions over fine- and coarse-grained variables, respectively. What they do have in common is a Dirac delta function that projects a fine-grained variable onto a coarse-grained observable. We now show that AF1 can be understood from the perspective of PK.
@@ 4.1 AF1’s potential as an instance of PK
In this section, we present a reformulation of the AF1 potential in probabilistic terms. First, we reinterpret the potential over the dihedral angles as an empirical prior that is updated with soft evidence on pairwise distances. Building on previous work concerning probabilistic models over dihedral angles [19, 24, 8, 39, 71, 40] based on directional statistics [46, 41, 51, 43, 44], this prior is formulated as
$\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta})$ $=\prod_{i}\textrm{vonMises}\left(\phi_{i};\boldsymbol{\Theta}_{i}\right)\textrm{vonMises}\left(\psi_{i};\boldsymbol{\Theta}_{i}\right),$ (17)
where the parameters are obtained from a deep model, $\boldsymbol{\Theta}=\textrm{NN}_{\textrm{Dihedral}}(\boldsymbol{s},\boldsymbol{M})$, and vonMises is the von Mises distribution. This circular distribution is analogous to a Gaussian distribution and suitable for modeling dihedral angles in $\pi,\pi] [$46]. The prior over dihedral angles implies a matching prior over distances,
$\pi(\boldsymbol{D};\boldsymbol{\varGamma})$ $\propto\prod_{i,j,i\neq j}p\left(D_{i,j};\boldsymbol{\varGamma}_{i,j}\right),$ (18)
where $\boldsymbol{\varGamma}=\textrm{NN}_{\textrm{Reference}}(L(\boldsymbol{s}),\mathbb{I}_{\textrm{Glycine}}(\boldsymbol{s}))$.
@@
$p(\boldsymbol{D};\boldsymbol{\Omega})$ $\propto\prod_{i,j,i\neq j}p\left(D_{i,j};\boldsymbol{\Omega}_{i,j}\right),$ (19)
where $\boldsymbol{\Omega}=\textrm{NN}_{\textrm{Distance}}(\boldsymbol{s},\boldsymbol{M})$. We now have all the elements in place to invoke the machinery of probability kinematics. Following PK’s RR update given by (14), the posterior is obtained as
$p(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta},\boldsymbol{\varGamma},\boldsymbol{\Omega})$ $\propto\frac{p(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\Omega})}{\pi(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\varGamma})}\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta}).$ (20)
By construction and following [8, 67], AF1 adopts the J-condition induced by a prior over dihedral angles. Why is the PK update needed, if dihedrals angles are in principle sufficient to determine the structure? The answer is well known: due to an elbow effect, the dihedral prior is accurate for residues that are close in sequence position, but becomes increasingly unreliable for residues that are distant [30, 23]. This locality is a well-known limitation of probabilistic models based on dihedral angles, and is due to the accumulation of small errors [4]. In AF1, PK provides a minimal and coherent update of the prior, incorporating soft evidence on Euclidean distances while harnessing the prior’s useful representation of local structure. The J-condition is illustrated for a concrete protein example in Appendix 10, and some practical aspects of AF1’s application of PK are discussed in Appendix 11.
AF1 predictions are obtained by gradient descent minimization in the space of the dihedral angles, $\boldsymbol{\phi},\boldsymbol{\psi}$ . This amounts to a maximum a posteriori (MAP) estimate obtained from the compositional empirical Bayes model constructed using PK. However, in contrast to PMF-based heuristics and point estimate prediction, PK allows the formulation of well-defined posterior distributions, as we demonstrate in the next section using a tractable synthetic model.
@@ 5 Experiments
Figure 3: Left: Empirical PDFs of the origin-to-end distance, $r$, for the unmodified VRW prior (with $\mu=0,\kappa=10$, $N=5$; blue histogram), the PK-updated VRW (red histogram), and a naive product of the $\textrm{ScaledBeta}(\alpha=10,\beta=10,N=5)$ target distribution with the VRW prior, i.e., without dividing by the reference distribution (grey histogram). The latter does not fit the target distribution, clearly showing that the reference distribution is required. Also shown are the PDFs of the target ScaledBeta distribution (black line) and of the theoretical distribution for $r$ (Stephens distribution; dashed line), showing an excellent fit in both cases. Right: 500 trajectories, $\boldsymbol{X}(\boldsymbol{\theta}$), sampled from the VRW prior (blue) and the PK-updated distribution (red). A dot denotes the end point of a random walk. As expected, the endpoints of the PK-updated distribution lie closer to the origin.
@@
To demonstrate PK in a way relevant to protein structure prediction, we formulate a synthetic model that by design has many crucial properties in common with the AF1 case. For both AF1 and the synthetic model, a generative model that specifies Euclidean coordinates based on an angular distribution is updated using distance evidence. Our experimental setup is deliberately minimal to demonstrate the feasibility and precision of PK updates in a tractable model, providing a bedrock for applications to more complex systems in future work, conditional on open solving problems such as formulating proper PDFs over distance matrices [15], which is beyond the scope of this article.
As prior, we use a von Mises random walk (VRW) in Euclidean 2D space. As coarse grained variable, we use the length of the resultant vector, that is, the distance between the origin and the end of the VRW, which has the advantage that its distribution is known [46, 63].
The similarities with AF1 will include the following elements: (a) both models make use of the von Mises distribution for a prior over angles, (b) the angles sampled from the von Mises distributions specify Euclidean coordinates that can be interpreted as atom positions, (c) subsequent atomic positions are at a given, fixed distance from each other, (d) the prior is updated with evidence concerning distances between a subset of the atoms, (e) the conditional PDF of angles given distances is not available, requiring the use of PK’s RR update given by (14), and (f) the updated atomic positions are less spatially extended than the prior (see Appendix 10).
@@
In this section, $\boldsymbol{X}$ will be an $(N+1)\times 2$ matrix of 2-dimensional coordinates, resulting from a VRW with $N$ steps. The VRW with mean $\mu$ and concentration (i.e., inverse variance) $\kappa$ is,
$\begin{cases}\boldsymbol{X}_{i}=(0,0),&i=0\\ \boldsymbol{X}_{i+1}=\boldsymbol{X}_{i}+\boldsymbol{v}(\theta_{i}),\theta_{i}\sim\textrm{vonMises}(\mu,\kappa),&0<i\leq N,\end{cases}$ (21)
@@
where $\boldsymbol{\theta}=\left\{\theta_{i}\right\}_{i=1}^{N}$ are angles in $\pi,\pi]$ and $\boldsymbol{v}(\theta_{i})$ is the corresponding 2D unit vector. The coordinates are fully determined by the angles, $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta})$. For this case, an approximate PDF of the Euclidean distance between start and end of the VRW, i.e., the length of the resultant vector, $r(\boldsymbol{X})=\left\|\boldsymbol{X}_{0}-\boldsymbol{X}_{N}\right\|{{}^{2}}$, is given by Stephens (1969),
$2N\gamma\left(1-\frac{r}{N}\right)\sim\chi_{N-1}^{2},$ (22)
@@
where $\frac{1}{\gamma}=\frac{1}{\kappa}+\frac{3}{8\kappa^{2}}$ [63]. For $\kappa\geqslant 4$, the approximation is excellent [63]. We will write $r\sim\textrm{Stephens}(r;\kappa,N).$ Note that the maximum value of $r$ is $N$, since each step has unit length. We define the prior distribution $\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa)$ as the generative model induced by the VRW. The marginal probability over $r$ for this prior can be written as
@@
$\pi_{\textrm{VRW}}(r;\kappa,N)=\textrm{Stephens}(r;\kappa,N)=\int\delta\left(r-r(\boldsymbol{\theta})\right)\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa,N)\textrm{d}\boldsymbol{\theta}.$ (23)
@@
Note that $r$ is a random variable, while $r(\boldsymbol{\theta})$ is the realization of $r$ for a specific VRW with angles $\boldsymbol{\theta}$. Now, we will demonstrate how PK allows us to update the prior using soft information on the distribution of $r$. We wish to modify the distribution over trajectories such that the marginal distribution of $r/N$ follows a $\textrm{Beta}(\alpha,\beta)$ distribution [^2] Note that as $r\in[0,N],$we divide by $N$ to be within the support of the Beta distribution. with given parameters $(\alpha,\beta)$ instead of $\textrm{Stephens}(r;\kappa)$. The target distribution over $r$ thus becomes after applying the change of variable formula,
$p(r;\alpha,\beta,N)=\frac{1}{N}\frac{1}{B(\alpha,\beta)}\left(\frac{r}{N}\right)^{\alpha-1}\left(1-\frac{r}{N}\right)^{\beta-1},\textrm{with }0<r<N,$ (24)

### v3

@@ 2.4 Probability kinematics
As in the classic example of Whitworth’s horses (Diaconis and Zabell, 1982) (see Appendix C), new information often does not arrive in the form of a deterministic event, but rather as a revision in the probability of various hypotheses. PK provides a framework for updating beliefs under such uncertain evidence. Unlike classical Bayesian conditioning, which updates a prior $\pi(\boldsymbol{\omega})$ to a posterior $p(\boldsymbol{\omega})\propto p(\boldsymbol{y}\mid\boldsymbol{\omega})\pi(\boldsymbol{\omega})$ using data $\boldsymbol{y}$, PK updating uses revised probabilities on a partition of the sample space of $\boldsymbol{\omega}$ (Figure 1).
@@
by construction making use of the Jeffrey- or J-condition $p(\boldsymbol{\omega}\mid\xi_{i})=\pi(\boldsymbol{\omega}\mid\xi_{i})$.
Unlike conventional Bayesian inference, PK does not rely on an explicit likelihood function, and its evidence is not a concrete event. Nonetheless, PK preserves essential features of Bayesian updating. First, PK satisfies the Dutch book argument (Armendt, 1980). Second, PK reduces to conventional Bayesian updating when for one of the partition elements, $p(\xi_{i})=1$. Third, among all distributions that match the updated marginals $\{p(\xi_{i})\}$, the PK update minimizes the Kullback-Leibler divergence to the prior $\pi$ (Diaconis and Zabell, 1982). This links the method to a variational principle (Blei et al., 2017; Kingma and Welling, 2019) that fits the Bayesian interpretation of probability (Diaconis and Zabell, 1982). PK is usually formulated in terms of probabilities and finite partitions (Meehan and Zhang, 2020). In the next section, we outline the more general probabilistic framework needed to address AF1.
@@ 3.1 PK for the case of PDFs and infinite partitions
We provide a definition tailored to the present case, based on (Diaconis and Zabell, 1982) and (Meehan and Zhang, 2020), while avoiding the formalism of measure theory and irrelevant boundary cases.
@@
  • A random variable $\boldsymbol{\omega}\in\Omega$, where $\Omega$ is a smooth manifold (e.g., the space of $m$ dihedral angles, the $m$-torus $\mathbb{T}^{m}$), equipped with a prior density $\pi(\boldsymbol{\omega})$.
  • A random variable $\boldsymbol{\xi}$, defined as a many-to-one function of $\boldsymbol{\omega}$. For $\boldsymbol{\omega}\in\Omega$, the corresponding realization is written $\boldsymbol{\xi}(\boldsymbol{\omega})\in\Xi$, where $\Xi$ is a smooth manifold (e.g., the space of valid pairwise distances between $n$ points in $\mathbb{R}^{3}$). The prior density of $\boldsymbol{\xi}$ is induced by that of $\boldsymbol{\omega}$ as
    $\pi(\boldsymbol{\xi})=\int\delta\!\big(\boldsymbol{\xi}(\boldsymbol{\omega})-\boldsymbol{\xi}\big)\,\pi(\boldsymbol{\omega})\,\mathrm{d}\boldsymbol{\omega}.$ (11)
    Each value $\boldsymbol{\xi}\in\Xi$ corresponds to the partition element $\{\boldsymbol{\omega}\in\Omega:\boldsymbol{\xi}(\boldsymbol{\omega})=\boldsymbol{\xi}\}$.
  • Evidence as a new density $p(\boldsymbol{\xi})$.
@@
    Probability kinematics update rule: The posterior $p(\boldsymbol{\omega})$ is obtained as
    $p(\boldsymbol{\omega})=\int\pi(\boldsymbol{\omega}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}.$ (14)
@@
    Reference ratio update rule: Equivalently, after applying Bayes’ formula to (14),
    $p(\boldsymbol{\omega})$ $=\int\frac{\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))\pi(\boldsymbol{\omega})}{\pi(\boldsymbol{\xi})}p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}$ (15)
    $=\frac{p(\boldsymbol{\xi}(\boldsymbol{\omega}))}{\pi(\boldsymbol{\xi}(\boldsymbol{\omega}))}\pi(\boldsymbol{\omega}).$ (16)
    The Dirac delta function arises from the fact that $p(\boldsymbol{\xi}\mid\boldsymbol{\omega})=\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))$. Note that the integrand is zero when $\boldsymbol{\xi}\neq\boldsymbol{\xi}(\boldsymbol{\omega})$, and that (16) follows directly from the J-condition given by (13). This formulation of PK is useful when the conditional in (14) is unavailable.
@@ 3.2 KBPs versus PMFs
KBPs can be understood as an (approximate) application of PK (Hamelryck et al., 2010; Mardia et al., 2011; Valentin et al., 2014; Hamelryck et al., 2015). Consider a prior distribution over protein structure, $\pi(\boldsymbol{X})$. Without any loss of generality, we leave out the conditioning on sequence. We are now given evidence $p(\boldsymbol{D})$ about pairwise distances between a subset of the coordinates, e.g., the $C\beta$ coordinates. Thus, the coarse-grained variable $\boldsymbol{D}$ induces a partition on the sample space of the fine-grained variable $\boldsymbol{X}$. These variables correspond to $\boldsymbol{\xi}$ and $\boldsymbol{\omega}$ in Definition 1, respectively. The PK update following (16) is then given by
$p(\boldsymbol{X})$ $=\int\frac{\delta(\boldsymbol{D}(\boldsymbol{X})-\boldsymbol{D})\pi(\boldsymbol{X})}{\pi(\boldsymbol{D})}p(\boldsymbol{D})\textrm{d}\boldsymbol{D}$ (17)
$=\frac{p(\boldsymbol{D}(\boldsymbol{X}))}{\pi(\boldsymbol{D}(\boldsymbol{X}))}\pi(\boldsymbol{X}).$ (18)
Note that $\boldsymbol{D}$ is a random variable, while $\boldsymbol{D}(\boldsymbol{X})$ is its realization at $\boldsymbol{X}$. A minus-log transform recovers the KBP given by (9),
@@
if we assume conditional independence of pairwise distances (see also Appendix E). The second term does not appear in standard KBP expressions, as it is typically implicitly accounted for by Monte Carlo sampling (Simons et al., 1997) or scoring of existing structures (Sippl, 1993b).
The difference between PK/KBPs and PMFs is now clear: while PMFs amount to a projection and marginalization of a given Boltzmann distribution of a single physical system as in (7), often formulated relative to a corresponding non-interacting system by subtracting a reference energy as in (8), PK/KBPs correspond to a generalized Bayesian update of a prior as in (14-16). Moreover, PK/KBPs and PMFs differ in producing distributions over fine- ($\boldsymbol{\omega}$) and coarse-grained ($\boldsymbol{\xi}$) variables, respectively, i.e.,
$\textrm{PMFs: }p(\boldsymbol{\xi})$ $=\int\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))p(\boldsymbol{\omega})\textrm{d}\boldsymbol{\omega}$ (20)
$\textrm{PK/KBPs: }p(\boldsymbol{\omega})$ $=\int\frac{\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))\pi(\boldsymbol{\omega})}{\pi(\boldsymbol{\xi})}p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}$ (21)
We now show that AF1’s potential can be understood from the perspective of PK.
@@ 4.1 AF1’s potential as an instance of PK
In this section, we present a reformulation of the AF1 potential in probabilistic terms. First, we reinterpret the potential over the dihedral angles as an empirical prior that is updated with soft evidence on pairwise distances. Building on previous work on probabilistic models of dihedral angles (Edgoose et al., 1998; Hamelryck et al., 2006; Boomsma et al., 2008; Lennox et al., 2009; Zhao et al., 2010; Lennox et al., 2010) based on directional statistics (Mardia and Jupp, 2009; Ley and Verdebout, 2017; Pewsey and García-Portugués, 2021; Mardia, 2025a, b), this prior is formulated as
$\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta})$ $=\prod_{i}\textrm{vM}\left(\phi_{i};\boldsymbol{\Theta}\right)\textrm{vM}\left(\psi_{i};\boldsymbol{\Theta}\right),$ (22)
where the parameters are obtained from a deep neural network (NN), $\boldsymbol{\Theta}=\textrm{NN}_{\textrm{Dihedral}}(\boldsymbol{s},\boldsymbol{M})$, and vM is the von Mises distribution. This circular distribution resembles a Gaussian distribution for large $\kappa$ and is suitable for modeling dihedral angles in $(-\pi,\pi]$ (Mardia and Jupp, 2009). The prior over dihedral angles implies a matching prior over distances, [^2] We discuss the incorrect assumption of independent pairwise distances in Appendix E.
$\pi(\boldsymbol{D};\boldsymbol{\Gamma})$ $\propto\prod_{i,j,i\neq j}p\left(D_{i,j};\boldsymbol{\Gamma}_{i,j}\right),$ (23)
where $\boldsymbol{\Gamma}=\textrm{NN}_{\textrm{Reference}}(L,\mathbb{I}_{\textrm{Gly}}(\boldsymbol{s}))$.
@@
$p(\boldsymbol{D};\boldsymbol{\Lambda})$ $\propto\prod_{i,j,i\neq j}p\left(D_{i,j};\boldsymbol{\Lambda}_{i,j}\right),$ (24)
where $\boldsymbol{\Lambda}=\textrm{NN}_{\textrm{Distance}}(\boldsymbol{s},\boldsymbol{M})$. We now have all the elements in place to invoke the machinery of PK. Following PK’s reference ratio update given by (16), which avoids the use of the unavailable conditional used in (14), the posterior is obtained as
$p(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta},\boldsymbol{\Gamma},\boldsymbol{\Lambda})$ $\propto\frac{p(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\Lambda})}{\pi(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\Gamma})}\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta}).$ (25)
By construction, AF1 adopts the J-condition induced by a prior over dihedral angles (Boomsma et al., 2008; Valentin et al., 2014). Why is the PK update needed, if dihedral angles are in principle sufficient to determine the structure? The answer is well known: due to an elbow effect, the dihedral prior is accurate for amino acids that are close together in the protein sequence, but becomes increasingly unreliable for amino acids that are distant (Holmes and Tsai, 2004; Gong et al., 2005). This locality is a well-known limitation of probabilistic models of proteins based on dihedral angles, and is due to the accumulation of small errors and improper covariance representation (Arts et al., 2024). In AF1, PK provides a minimal and coherent update of the prior, incorporating soft evidence on Euclidean distances while harnessing the prior’s useful representation of local structure. The J-condition adopted by AF1 is further discussed in Appendix D. Some practical, approximate aspects of AF1’s application of PK, including modeling the pairwise distances as conditionally independent and approximating the reference distribution, are discussed in Appendix E.
AF1 predictions are obtained by gradient descent minimization in the space of the dihedral angles, $\boldsymbol{\phi},\boldsymbol{\psi}$. This approximates a maximum a posteriori (MAP) estimate obtained from the compositional empirical Bayes model constructed using PK. However, PK allows the formulation of well-defined posterior distributions involving angles, coordinates and distances, as we demonstrate in the next section using a tractable synthetic model.
@@ 5 EXPERIMENTS
Figure 3: (left) Empirical PDFs of the origin-to-end distance, $d$, for the unmodified VRW prior (with $\mu=0,\kappa=10$, $N=5$; blue histogram), the PK-updated VRW (red histogram), and a naive product of the $\textrm{ScaledBeta}(\alpha=10,\beta=10,N=5)$ target distribution with the VRW prior, i.e., without dividing by the reference distribution (gray histogram). The latter does not fit the target distribution, clearly showing that the reference distribution is required. Also shown are the PDFs of the target ScaledBeta distribution (black line) and of the theoretical distribution for $d$ (Stephens distribution; dashed line), showing an excellent fit in both cases. (right) 500 2D trajectories, $\boldsymbol{\chi}(\boldsymbol{\theta})$, sampled from the VRW prior (blue) and the PK-updated distribution (red). A dot denotes the end point of a random walk. As expected, the endpoints of the PK-updated distribution lie closer to the origin.
@@
To demonstrate PK in a way relevant to protein structure prediction, we formulate a synthetic model that by design has many crucial properties in common with the AF1 case (summarized in Appendix B). For both AF1 and the synthetic model, a generative model that specifies Euclidean coordinates based on an angular distribution is updated using distance evidence.
Our synthetic model is deliberately minimal to demonstrate the feasibility and precision of PK updates involving angles, coordinates and distances in a tractable model [^3] We provide an application to a small 3D peptide model, which requires estimation of the reference distribution, in Appendix F.. Our probabilistic framework and synthetic model are thus meant to provide a bedrock for applications to more complex systems in future work.
As prior, we use a von Mises random walk (VRW) in Euclidean 2D space. As coarse grained variable, we use the length of the resultant vector, that is, the distance between the origin and the end of the VRW, which has the advantage that its distribution is known (Mardia and Jupp, 2009; Stephens, 1969).
The similarities with AF1 include the following elements: (a) both models make use of the von Mises distribution for a prior over angles, (b) the angles sampled from the von Mises distributions specify Euclidean coordinates that can be interpreted as atom positions, (c) subsequent atomic positions are at a given, fixed distance from each other, (d) the prior is updated with evidence concerning distances between a subset of the atoms, (e) the conditional PDF of angles given distances is not available, requiring the use of PK’s reference ratio update given by (16), and (f) the updated atomic positions are less spatially extended than the prior (see also Appendix D and Figure 5). Specifically, as we detail below, angle vector $\boldsymbol{\theta}$, 2D coordinates $\boldsymbol{\chi}$ and distance $d$ of the synthetic model correspond to $(\boldsymbol{\phi},\boldsymbol{\psi}),\boldsymbol{X}$ and $\boldsymbol{D}$ of AF1, respectively. The correspondences between AF1 and the synthetic model are summarized in Appendix B.
@@
Let $\boldsymbol{\chi}$ be an $(N+1)\times 2$ matrix of 2-dimensional coordinates, resulting from a VRW with $N$ steps. The VRW with mean $\mu$ and concentration (i.e., inverse variance) $\kappa$ is,
$\begin{cases}\boldsymbol{\chi}_{i}=(0,0),&i=0\\ \boldsymbol{\chi}_{i}=\boldsymbol{\chi}_{i-1}+\boldsymbol{v}(\theta_{i}),&0<i\leq N,\end{cases}$ (26)
@@
where $\boldsymbol{\theta}=\left\{\theta_{i}\sim\textrm{vM}(\mu,\kappa)\right\}_{i=1}^{N}$ are angles in $(-\pi,\pi]$ and $\boldsymbol{v}(\theta_{i})$ is the corresponding 2D unit vector. As in AF1, the coordinates are fully determined by the angles; as before, we will write $\boldsymbol{\chi}(\boldsymbol{\theta})$ for the realization of $\boldsymbol{\chi}$ at $\boldsymbol{\theta}$. For this case, an approximate PDF of the Euclidean distance between start and end of the VRW, i.e., the length of the resultant vector, $d(\boldsymbol{\chi})=\left\|\boldsymbol{\chi}_{0}-\boldsymbol{\chi}_{N}\right\|$, is given by Stephens (1969),
$2N\gamma\left(1-\frac{d}{N}\right)\sim\textrm{Chi}_{N-1}^{2},$ (27)
@@
where $\frac{1}{\gamma}=\frac{1}{\kappa}+\frac{3}{8\kappa^{2}}$. For $\kappa\geqslant 4$, the approximation is excellent. We will write $d\sim\textrm{Stephens}(d;\kappa,N).$ Note that the maximum value of $d$ is $N$, since each step has unit length. We define the prior distribution $\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa)$ as the generative model induced by the VRW. The marginal probability over $d$ for this prior is
@@
$\pi_{\textrm{VRW}}(d;\kappa,N)=\textrm{Stephens}(d;\kappa,N)\\ =\int\delta\left(d-d(\boldsymbol{\theta})\right)\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa,N)\textrm{d}\boldsymbol{\theta}.$ (28)
@@
Note that $d$ is a random variable, while $d(\boldsymbol{\theta})$ is the realization of $d$ for a specific VRW with angles $\boldsymbol{\theta}$. Now, we will demonstrate how PK allows us to update the prior using soft information on the distribution of $d$. We wish to modify the distribution over trajectories such that the marginal distribution of $d/N$ follows a $\textrm{Beta}(\alpha,\beta)$ distribution [^4] Note that as $d\in[0,N],$we divide by $N$ to be within the support of the Beta distribution. with given parameters $(\alpha,\beta)$ instead of $\textrm{Stephens}(d;\kappa)$. The target distribution over $d$ thus becomes after applying the change of variable formula,
$p(d;\alpha,\beta,N)=\\ \frac{1}{N}\frac{1}{B(\alpha,\beta)}\left(\frac{d}{N}\right)^{\alpha-1}\left(1-\frac{d}{N}\right)^{\beta-1},$ (29)
where $B(\alpha,\beta)$ is the standard normalization constant of the Beta distribution and $d\in[0,N]$. We will write $d\sim\textrm{ScaledBeta}(d;\alpha,\beta,N)$.